Jen invested $7,500 in two bank accounts, one paying 4% simple interest, and the other paying 6%. If her annual interest from the two accounts is $420, how much did she invest at 6%?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
7500
Interest----- 420.00

Part I 4.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 7500 -------------1
Interest on both investments
4.00% x + 6.00% y= 420
Multiply by 100
4 x + 6 y= 42000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -30000.00
Add this to (2)
0 x 2 y= 12000
divide by 2
y = 6000 - x)
Part I 4.00% $ 1500
Part II 6.00% $ 6000

CHECK
1500 --------- 4.00% ------- 60.00
6000 ------------- 6.00% ------- 360.00
Total -------------------- 420.00

m.ananth@hotmail.ca